40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Dec 21 18:18:03 2020, gautam, Update, CDS, Many model changes CDSoverview.png
    Reply  Tue Dec 22 22:11:37 2020, gautam, Update, CDS, AA filt repaired and reinstalled 
Message ID: 15743     Entry time: Mon Dec 21 18:18:03 2020     Reply to this: 15744
Author: gautam 
Type: Update 
Category: CDS 
Subject: Many model changes 

The CDS model change required to get the AS WFS signals into the RTCDS system are rather invasive.

  • We use VCS for these models. Linus Torvald may question my taste but I also made local backups of the models, just in case...
  • Particularly, the ADC1 card on c1ioo is completely reconfigured.
  • I also think it's more natural to do all the ASC computations on this machine rather than the c1lsc machine (it is currently done in the c1ass model). So there are many IPC changes as well.
  • I have documented everything carefully, and the compile/install went smoothly.
  • Taking down all the FE servers at 1830 local time
    1. To propagagte the model changes
    2. To make a hardware change in the c1rfm card in the c1ioo machine to configure it as "ROGUE MASTER 0"
    3. To clear the RFM errors we are currently suffering from will require a model reboot anyways.
  • Recovery was completed by 1930 - the RFM errors are also cleared, and now we have a "ROGUE MASTER 👾" on the network. Pretty smooth, no major issues with the CDS part of the procedure to report.
  • The main issue is that in the AA chassis I built, Ch14 (with the first channel as Ch1) has the output saturated to 28V (differential). I'm not sure what kind of overvoltage protection the ADC has - we frequently have the inputs exceed the spec'd +/-20 V (e.g. when the whitening filters are engaged and the cavity is fringing), but pending further investigation, I am removing the SCSI connection from the rear of the AA chassis.

In terms of computational load, the c1ioo model seems to be able to handle the extra load no issues - ~35us/60us per cycle. The RFM model shows no extra computational time

After this work, the IMC locking and POX/POY locking, and dither alignment servos are working okay. So I have some confidence that my invasive work hasn't completely destroyed everything. There is some hardware around the rear of 1X2 that I will clear tomorrow.

Attachment 1: CDSoverview.png  31 kB  Uploaded Mon Dec 21 19:50:42 2020  | Hide | Hide all
CDSoverview.png
ELOG V3.1.3-